• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fansler

    Log entry time 10:30:16 on March 21, 2014

    Entry number 387102

    keyword=LD1

    The left D1 was displaying symptoms of a heat load. We checked all associated vacuums and discovered the magnet at 14 mtorr. There is a blower attached to the magnet, we should be able to remove it on Monday. The bogies are locked out, the blower will restrict movement of the left spectrometer. If the left spectrometer needs to be moved over the weekend, contact the tech on call.



    A copy of this log entry has been emailed to: folts@jlab.org, glorioso@jlab.org